OpenStudy (anonymous):

Find the slope of the line. 4x + 5y = -2

OpenStudy (anonymous):

-4/5

OpenStudy (anonymous):

\[4x+5y=-2\] 5y=-4x-2 y=-4x/5-2/5 From this the slope is like esamalaa said -4/5

OpenStudy (anonymous):

Rule 1: Convert this equation to the form: y=mx + b

OpenStudy (anonymous):

m=the slope of the line